OpenStudy (anonymous):

quadratic formula. 2x^2-14=0

OpenStudy (crashonce):

ok so first factor out the two: 2x(x^2-7) SOlve from there

OpenStudy (yttrium):

Use traditional way of solving this. \[2x^2 - 14 = 0\] \[2x^2 = 14\] \[x^2 = 7\] hence, \[x = \pm \sqrt7\]

OpenStudy (crashonce):

you could dot that thats better

OpenStudy (anonymous):

okay thank you both. will you help me with one more?

OpenStudy (crashonce):

sure

OpenStudy (anonymous):

g^2+3g=0

OpenStudy (crashonce):

well make it : g^2=-3g then divide both sides by g: g=-3

OpenStudy (anonymous):

thank you!

OpenStudy (crashonce):

no probs